OpenStudy (anonymous):

At a canning company the daily production cost, y, is given by the quadratic equation y = 650 – 15x + 0.45x2, where x is the number of canned items. What is the MINIMUM daily production cost?

OpenStudy (stamp):

max/min of a parabola ax^2+bx+c is described by\[\frac{-b}{2a}\]

OpenStudy (anonymous):

what is the answer then?

OpenStudy (stamp):

0.45xx-15x+650 a = 0.45 b = -15 -b/(2a) = -(-15)/(0.90) 15 / 0.9 = 15 * 10/9 = 150 / 9

OpenStudy (anonymous):

A) $1,025.00 B) $1,186.25 Eliminate C) $525.00 D) $536.25

OpenStudy (anonymous):

is it A, B, C, or D

OpenStudy (anonymous):

there is no f what do you mean

OpenStudy (stamp):

Plug in x = 150/9

OpenStudy (stamp):

The production cost is y. The minimum y value occurs at x = 150/9. So to find the answer to the question, the minimum production value, you plug x = 150/9 and find y = answer

OpenStudy (anonymous):

i did and i didnt get anywhere close to the answer

OpenStudy (anonymous):

how about this last one.... A ball is kicked upward with an initial velocity of 56 feet per second. The ball's height h (in feet) can be expressed as a function of time t (in seconds) by the equation h = -16t2 + 56t. How much time does the ball take to return to the ground?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can you please help me